It was our first LBTI night of 2013, last night. The goal was to do phase sensing engineering the first half and exoplanet hunting (LEECH) the second half.
We opened at sunset and closed the right side AO. Left side was causing more problems, but finally fixed the alignment to get it closed. We scanned for fringes and found them quickly. However, seeing was 1.5-2". The phase sensor was aligned and run at 400-500 Hz. However, fringes were not detectable in the poor seeing.
Switched over to LMIRCam at 12:30, only to have first the winds come up and then the humidity, requiring us to close for the rest of the night.
/groups/lbti/search/index.rss?tag=hotlist/groups/lbti/search/?tag=hotWhat’s HotHotListHot!?tag=hot1/groups/lbti/sidebar/HotListAdded tag - hotphinzPhil Hinz2010-10-15 20:37:53+00:002010-10-15 20:37:53addTag3phinzPhil Hinz2010-10-15 13:18:40+00:002010-10-15 13:18:40updated2First createdphinzPhil Hinz2010-10-15 13:05:39+00:002010-10-15 13:05:39created1weblog2010-10-15T13:18:40+00:00groups/lbti/weblog/b808eFalseFirst Night Results/groups/lbti/weblog/b808e/First_Night_Results.htmlPhil Hinz3 updatesFirst Night Results
It was an eventful first night on the LBT. Network connections failed, hardware refused to turn on, all the typical problems of a new instrument. Aft...Falsephinz2010-10-15T13:18:40+00:00hot/groups/lbti/search/index.rss?sort=modifiedDate&kind=all&sortDirection=reverse&excludePages=wiki/welcomelist/groups/lbti/search/?sort=modifiedDate&kind=all&sortDirection=reverse&excludePages=wiki/welcomeRecent ChangesRecentChangesListUpdates?sort=modifiedDate&kind=all&sortDirection=reverse&excludePages=wiki/welcome0/groups/lbti/sidebar/RecentChangesListmodifiedDateallRecent ChangesRecentChangesListUpdateswiki/welcomeNo recent changes.reverse5searchlist/groups/lbti/calendar/Upcoming EventsUpcomingEventsListEvents1Getting events…